OpenStudy (anonymous):

2. Simplify (2t^–3)^3(0.4r)^2 Thanks

OpenStudy (anonymous):

You have:\[(2t^{–3})^3(0.4r)^2 \]So this is:\[2^3(t^{-3})^30.4^2r^2\] Of course, 2³=8 and (0.4)²=0.16. 8*0.16=1.28, so that's fine. I see two problems in your answer: 1. t^-6 remember: (a^b)^c=a^(b*c), so... 2. "over" if you want to make a fraction with t in the denominator, you'll lose the - of the exponent.

OpenStudy (anonymous):

my answer choices are 3.2r^2 1.28r^2/t^-6 1.28r^2/t^9 0.8r^2/t^-6

OpenStudy (anonymous):

3.2r^2